OpenStudy (anonymous):

Please help!!!!!!! A river flows at the rate of 2mph. If a barge travels 3 miles upstream and back in two hours, what is the speed of the barge in still water?

OpenStudy (kropot72):

Let s be the speed of the barge in still water. Then the speed upstream is (s - 2) mph and the speed downstream is (s + 2) mph. \[time=\frac{distance}{speed}\] Time to travel upstream is \[\frac{3}{s-2}\ hours\] Time to travel downstream is \[\frac{3}{s+2}\] Total time to travel upstream 3 miles and return is 2 hours. Therefore \[2=\frac{3}{s-2}+\frac{3}{s+2}\] \[2=\frac{3(s+2)+3(s-2)}{s ^{2}-4}..............(1)\] Equation (1) can be rearranged to give the quadratic \[2s ^{2}-6s-8=0...........(2)\] Now solve the quadratic to find s.